OpenStudy (iwanttogotostanford):

HELP W/ ACT PREP QUESTION: Which of the following equations expresses c in terms of a for all real numbers, a, b, and c such that a^3=b and b^2=c? 1.) c=a^6 2.) c= a^5 3.) c=2a^3 4.) c=1/2 a 5.) c= a

OpenStudy (iwanttogotostanford):

@Vocaloid @zepdrix @pooja195

OpenStudy (iwanttogotostanford):

@Nnesha

jimthompson5910 (jim_thompson5910):

replace b with a^3 b^2 = c (b)^2 = c (a^3)^2 = c I'll let you finish up

OpenStudy (iwanttogotostanford):

@jim_thompson5910 im not sure how to finish! Also, what was your reassigning in knowing how to do that??

jimthompson5910 (jim_thompson5910):

I replaced b with a^3 because the first equation is a^3 = b or b = a^3 a^3 is the same as b, which is why the replacement can be done

jimthompson5910 (jim_thompson5910):

you'll use \[\LARGE (x^a)^b = x^{a*b}\]

jimthompson5910 (jim_thompson5910):

eg: \[\LARGE (x^8)^5=x^{8*5}=x^{40}\]

OpenStudy (iwanttogotostanford):

is that ALg. 2??

jimthompson5910 (jim_thompson5910):

yeah you'd most likely learn that in alg 2

jimthompson5910 (jim_thompson5910):

maybe alg 1, it depends on the teacher

OpenStudy (iwanttogotostanford):

so it would be choice number 1?

OpenStudy (iwanttogotostanford):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, c = a^6